A pro basketball player has a vertical leap of 25 inches, what is his hang time? Use the hang time function V=48T^2
Answer by nerdybill(7384)

A pro basketball player has a vertical leap of 25 inches, what is his hang time? Use the hang time function V=48T^2
.
just set V to 25 and solve for t:
25 = 48t^2
25/48 = t^2
sqrt(25/48) = t
0.72 secs = t